OpenStudy (anonymous):

Ivar bought a house for $375,900. He obtained a mortgage for $332,450. The documentary stamp tax fee for mortgages in his state is $0.35 per $100 or portion thereof. How much did he pay for the documentary stamp tax?

OpenStudy (anonymous):

1163.78??? idk what i just did :P

OpenStudy (anonymous):

\[.35\times \frac{332,450}{100}=.0035\times 332,450\]

OpenStudy (anonymous):

then a calculator

OpenStudy (anonymous):

looks like you get the answer above

OpenStudy (anonymous):

yay! :P